Which of the following values of x will make the sequence 3x, 4x+2, 7x-1 a geometric sequence?
1
Expert's answer
2013-01-29T08:27:14-0500
if 3x, 4x+2, 7x-1 is geometric sequence then
3x/(4x+2)=(4x+2)/(7x-1)3x(7x-1)=(4x+2)^2
21x^2-3x=16x^2+16x+4
5x^2-19x-4=0
D=19^2+4*5*4=361+80=441=(21)^2
so x1,2=( 19+-21)/10
so x1=4 and x2=-0.2
so only for these values we get geometric sequence